Water Movers, Inc. is known throughout the world for its H2O-X high-capacity water pump, used in irrigation systems. The pump’s standard cost is as follows. The company’s predetermined fixed overhead rate is based on an expected capacity of 100,000 direct labor hours per month.

Direct materials Direct labor Variable overhead Fixed overhead Standard Price Standard Quantity Standard Cost $96 40 36 20 $192 $6 per pound $10 per DLH $9 per DLH $5 per DLH 16 pounds 4 DLH 4 DLH 4 DLH During the month of September, the company produced 22,210 of the 25,000 pumps that had been scheduled for production in the budget. The company used 406,600 pounds of material during September. The direct labor payroll for the month was $1,012,000 for 92,000 direct labor hours. Variable overhead costs were $813,000; fixed overhead costs were $537,700. The companys purchasing agent signed a new supply contract that resulted in purchases of 526,300 pounds of direct materials at a total price of $2,894,650 Calculate Water Movers direct materials, direct labor, and overhead variances for September. (Round per unit values to 2 decimal places, e.g. 15.25 & final answers to 0 decimal places, e.g. 5,725. If variance is zero, select Not Applicable and enter o for the amounts.)
Answer #1
Standard Actual
SQ / SH SR Total AQ / AH AR Total
Direct Material    355,360.00           6.00      2,132,160.00    406,600.00           5.50    2,236,300.00
Direct Labor      88,840.00         10.00          888,400.00      92,000.00         11.00    1,012,000.00
Variable Overhead      88,840.00           9.00          799,560.00      92,000.00           8.84        813,000.00
Actual Production = 22,210 Pumps
Direct Material Price Variance = (SR - AR) X AQ Purchased
Direct Material Price Variance = ($6 - $5.50) X 526,300 pounds
Direct Material Price Variance = $263,150 (F)
Direct Materials Quantity Variance = (SQ - AQ) X SR
Direct Materials Quantity Variance = (355,360 pounds - 406,600 pounds) X $6
Direct Materials Quantity Variance = $307,440 (U)
Direct Labor Rate Variance = (SR - AR) X AH
Direct Labor Rate Variance = ($10 - $11) x 92,000 Hrs
Direct Labor Rate Variance = $92,000 (U)
Direct Labor Efficiency Variance = (SH - AH) X SR
Direct Labor Efficiency Variance = (88,840 hrs - 92,000 hrs) X $10
Direct Labor Efficiency Variance = $31,600 (U)
Variable Overhead Spending Variance = (SR - AR) X AH
Variable Overhead Spending Variance = ($9 -$8.84) X 92,000 hrs
Variable Overhead Spending Variance = $15,000 (F)
Variable Overhead Efficiency Variance = (SH - AH) X SR
Variable Overhead Efficiency Variance = (88,840 hrs - 92,000 hrs) X $9
Variable Overhead Efficiency Variance = $28,440 (U)
Fixed Overhead Spending Variance = Budgeted Overhead - Actual Overhead
Fixed Overhead Spending Variance = $500,000 - $537,700
Fixed Overhead Spending Variance = $37,700 (U)